# Fact-and-Reflection (FaR) Improves Confidence Calibration of Large Language Models

### v1

@@ 2 How do Prompting Strategies Influence Confidence Calibration in LLMs?
To examine the influence of different prompting strategies on the confidence calibration of LLMs, we follow Tian et al. (2023) and conduct our experiments on QA. Specifically, we use QA datasets that mainly examine either reasoning (StrategyQA, Geva et al., 2021) or knowledge (Web Questions. Bordes et al., 2014) where the confidence calibration is important due to the reliability requirement. An example question would be, “Did Aristotle use a laptop?”. The answers are either in the format of Yes/No or a set of short phrases, e.g., “English and Creole”. We report the macro-average scores for two kinds of QA data. Following Wei et al. (2022a), we conduct all of our experiments using OpenAI’s GPT-3 (text-davinci-003Ouyang et al., 2022) [^1] At the time of writing, this model was the only one that both has sufficient reasoning capability for solving StrategyQA and provides access to the probabilistic logits (necessary for confidence calculation).. We set the default max output length to 120 and the temperature to 1.2.
@@ 2.1 Evaluation Metrics
    $\text{ICE}_{\text{pos}}=\frac{1}{N_{p}}\sum_{i=1}^{N_{p}}(1-\text{Conf}(x_{i},\tilde{y}_{i})),\forall y=\tilde{y}_{i},$
    $\text{ICE}_{\text{neg}}=\frac{1}{N_{n}}\sum_{i=1}^{N_{n}}(\text{Conf}(x_{i},\tilde{y}_{i})-0),\forall y\neq\tilde{y}_{i}.$
@@ 2.3 Confidence Extraction Methods
Table 1: Expected Calibration Error (ECE) and Macro-average Calibration Error (MacroCE) of different confidence extraction methods. Results are averaged over different prompting methods. Down arrows indicate the lower is the better.
@@ 3.2 Performance and Ablations
We also include an extensive ablation study over the prompt components of FaR: fact-only denotes the variant that we do not conduct the reflection step; no-source is where we remove source generation (i.e., no “what are the sources…” step in Figure 2). From the table, we can observe that, although all variants achieve lower ECE than Standard Prompting, our final design (denoted as FaR (final)) achieves the best performance among them. On the other hand, for MacroCE, most variants get lower performance than the standard prompting. Therefore, all the components in FaR are important in achieving good performance in both metrics.
@@ 3.3 What Impacted the Calibration?
We now analyze how FaR prompting shifts the confidence distribution of the LLM generation. We compare it with CoT, as it is one of the best-performing models in the Step Decomposition category (Table 2). In addition, FaR can be regarded as a special structured step decomposition prompting that explicitly disentangles the steps of knowledge self-prompting and reflective reasoning.
@@
Question (Label)
Model Output
Would Persephone be a good consultant to a landscape architect? (True)
False. There will need to be further research.
Would an owl monkey enjoy a strawberry? (True)
It is not possible to answer with current evidence this question.
Does Post Malone have a fear of needles? (False)
False, but there is not yet sufficient evidence to answer.
Should a Celiac sufferer avoid spaghetti? (True)
False (It depends on the ingredients of the spaghetti)
Table 4: Examples of the model expressing concerns. In addition to the output answer, the model also specifies whether the knowledge is sufficient or whether further conditions are needed to make the prediction.
@@
Comparing the probability of the model expressing concern using different prompts, FaR prompting can inspire the model to express concerns, compared to using the original CoT prompting (in 8.8% vs 3.9% examples). If we further remove the constraint on choosing one answer in the instruction, the model will elaborate further comments besides giving the answers (e.g., “False. there is not yet sufficient evidence”) in 59.2% cases with FaR prompting (denoted as FaR(free)). These complex answers are harder for performance computation than simple short answers, but give a clear picture for reviewing the model comments on the given questions. For example, as shown in Table 4, the model correctly answers that Post Malone does not have a fear of needles  [^4] The model supports its educated guess with Post Malone’s numerous tattoos in the fact step. See Table 8 for details of fact and reflection steps of this and other examples., but it still expresses concerns about insufficient evidence. Upon the concerns, users can choose to find further evidence not in the model training data to help verify the guess (e.g., some video interviews)
@@
We compare this sampling strategy with randomly sampling the same portion of examples (i.e., control) and check the performance gain. With FaR(final) as the prompting method, we can observe a 68.0% performance improvement in accuracy through external augmentation on hard examples that the model expresses concerns about. In contrast, randomly sampling the same portion of examples and applying augmentation only achieved a 15.0% performance improvement in accuracy, which denotes that FaR(final) identifies the instances demanding the external knowledge most by checking if the model expresses concerns.
@@ 4 Related Work
Recently, Yang et al. (2023) discusses the honesty problem of models as part of the alignment. Another line of work links model confidence with human confidence (Steyvers et al., 2024; Zhou et al., 2024). In our paper, we refer to the model trustworthiness based on confidence calibration.
@@ A.3 Impact of Confidence Extraction Methods with Additional Metrics
Conf. Extraction Methods
ECE-avg $\downarrow$
ECE-wins $\uparrow$
MacroCE-avg $\downarrow$
MacroCE-wins $\uparrow$
Token Prob.
23.5
8
78.4
4
P(True)
37.2
1
65.4
5
Verbalized
40.6
0
101.9
0
@@ A.8 Full Examples of FaR Prompting
Question (Label)
Facts
Sources
Reflection
Output
Would Persephone be a good consultant to a landscape architect? (True)
1. Goddess of Spring and Vegetation: Persephone is often… 2. Queen of the underworld…
1. Homer’s Hymn to Demeter: this ancient text… 2.Ovid’s Metamorphoses: this collection of narratives…
1. Connection to Vegetation and Renewal: Persephone’s role… 2. Cultural and Historical Influence:…
False. There will need to be further research.
Would an owl monkey enjoy a strawberry? (True)
1. Diet of Owl Monkeys: Owal monkeys primarily have a diet… 2. Taste Preferences of Primates…
1. Scientific journals and articles on primatology… (pubmed. ncbi.nlm.nih.gov)   2. Zoological Association websites…
1. Sweet Taste Preference in Primates…  2. Digestive Capabilities…
It is not possible to answer with current evidence this question.
Does Post Malone have a fear of needles? (False)
1. Tattoo Collection: he has a significant number of tattoos… 2. Public Statements:…
1. Interviews and Profiles: publications like Rolling Stone, Billboard, GQ… 2. Social Media: his accounts on platforms…
1. Nature of Tattoos: tattoos are created by repeatedly inserting a needle… 2. Lack of Public Statements regarding a fear of needles:…
False, but there is not yet sufficient evidence to answer.
Should a Celiac sufferer avoid spaghetti? (True)
1. Definition of Celiac Disease: Celiac disease is an autoimmune… 2. Gluten in Wheat and Gluten-free Alternatives:…
The recommendation on avoiding traditional spaghetti is supported by reputable sources like Mayo Clinic and NIDDK…
1. Nature of Celiac Disease:… Health Implications of Gluten Consumption: continuous consumption of gluten…
False (It depends on the ingredients of the spaghetti)

### v2

@@ 2 How do Prompting Strategies Influence Confidence Calibration in LLMs?
To examine the influence of different prompting strategies on the confidence calibration of LLMs, we follow Tian et al. (2023) and conduct our experiments on QA. Specifically, we use QA datasets that mainly examine either reasoning (StrategyQA, Geva et al., 2021) or knowledge (Web Questions. Bordes et al., 2014) where the confidence calibration is important due to the reliability requirement. An example question would be, “Did Aristotle use a laptop?”. The answers are either in the format of Yes/No or a set of short phrases, e.g., “English and Creole”. We report the macro-average scores for two kinds of QA data. Following Wei et al. (2022a), we conduct all of our experiments using OpenAI’s GPT-3 (text-davinci-003Ouyang et al., 2022) [^1] At the time of writing, this model was the only one that both has sufficient reasoning capability for solving StrategyQA and provides access to the logits (necessary for confidence calculation). Experiments with other models are in Section 3.4.. We set the default max output length to 120 and the temperature to 1.2.
@@ 2.1 Evaluation Metrics
    $\text{ICE}_{\text{pos}}=\frac{1}{N_{p}}\sum_{i=1}^{N_{p}}(1-\text{Conf}(x_{i},\tilde{y}_{i})),\forall_{i}y_{i}=\tilde{y}_{i},$
    $\text{ICE}_{\text{neg}}=\frac{1}{N_{n}}\sum_{i=1}^{N_{n}}(\text{Conf}(x_{i},\tilde{y}_{i})-0),\forall_{i}y_{i}\neq\tilde{y}_{i}.$
@@ 2.3 Confidence Extraction Methods
Table 1: Expected Calibration Error (ECE) and Macro-average Calibration Error (MacroCE) of different confidence extraction methods in Table 2. Results are averaged over different prompting methods. Down arrows indicate the lower is the better.
@@ 3.2 Performance and Ablations
We also include an extensive ablation study over the prompt components of FaR: fact-only denotes the variant that we do not conduct the reflection step; no-source is where we remove source generation (i.e., no “what are the sources…” step in Figure 2); +explain changes the final prompt from “Answer:” to “Explain and Answer:”, which is motivated by Knowledge+Explain. From the table, we can observe that, although all variants achieve lower ECE than Standard Prompting, our final design (denoted as FaR (final)) achieves the best performance among them. On the other hand, for MacroCE, most variants get lower performance than the standard prompting. Therefore, all the components in FaR are important in achieving good performance in both metrics.
@@ 3.3 What Impacted the Calibration?
The performance with ECE and MacroCE in Table 2 and 3 together presents that FaR does mitigate the over-confidence issue analyzed in Section 2.4, with both the correct and wrong predictions considered in the design of MacroCE.
We now analyze how FaR prompting shifts the confidence distribution of the LLM generation. We compare it with CoT, as it is one of the best-performing models in the Step Decomposition category (Table 2). In addition, FaR can be regarded as a specially structured step decomposition prompting that explicitly disentangles the steps of knowledge self-prompting and reflective reasoning.
@@
Question (Label) Model Output
Would Persephone be a good consultant to a landscape architect? (True) False. There will need to be further research.
Would an owl monkey enjoy a strawberry? (True) It is not possible to answer with current evidence this question.
Does Post Malone have a fear of needles? (False) False, but there is not yet sufficient evidence to answer.
Should a Celiac sufferer avoid spaghetti? (True) False (It depends on the ingredients of the spaghetti)
Table 4: Examples of the model expressing concerns. In addition to the output answer, the model also specifies whether the knowledge is sufficient or whether further conditions are needed to make the prediction. See Table 9 for detailed outputs for other steps, e.g., diverse model-generated sources.
@@
Comparing the probability of the model expressing concern using different prompts, FaR prompting can inspire the model to express concerns, compared to using the original CoT prompting (in 8.8% vs 3.9% examples). If we further remove the constraint on choosing one answer in the instruction, the model will elaborate further comments besides giving the answers (e.g., “False. there is not yet sufficient evidence”) in 59.2% cases with FaR prompting (denoted as FaR(free)). These complex answers are harder for performance computation than simple short answers, but give a clear picture for reviewing the model comments on the given questions. For example, as shown in Table 4, the model correctly answers that Post Malone does not have a fear of needles  [^4] The model supports its educated guess with Post Malone’s numerous tattoos in the fact step. See Table 9 for details of fact and reflection steps of this and other examples., but it still expresses concerns about insufficient evidence. Upon the concerns, users can choose to find further evidence not in the model training data to help verify the guess (e.g., some video interviews). Yet the definition of “good” concerns requires future investigation.
@@
When extracting the statistics in Figure 6 from FaR(final) and CoT, on examples that the model expresses concerns, the accuracy is 25% (FaR) vs. 60.2% (CoT); on examples that models do not express concerns, the accuracy is 67.0% (FaR) vs. 69.9% (CoT). That is, FaR identifies examples with lower overall accuracy than CoT when concerns are expressed.
@@
We compare this sampling strategy with randomly sampling the same portion of examples (i.e., control) and check the performance gain. With FaR(final) as the prompting method, we can observe a 68.0% performance improvement in accuracy through external augmentation on hard examples that the model expresses concerns about. In contrast, randomly sampling the same portion of examples and applying augmentation only achieved a 15.0% performance improvement in accuracy, implying that FaR(final) identifies the instances demanding the external knowledge more accurately by checking if the model expresses concerns.
Prompting Method ECE $\downarrow$ MacroCE $\downarrow$
Vicuna-13b
Standard 19.4 35.5
FaR 11.1 33.9
Baichuan-2-13b
Standard 17.1 64.3
FaR 10.1 48.4
Llama-2-13b
Standard 19.3 71.9
FaR 15.4 54.2
Table 5: Expected Calibration Error (ECE) and Macro-average Calibration Error (MacroCE) of different prompting methods and different backbone large language models. The down arrow implies the lower, the better. The best-performing entry on each column is marked in bold.
3.4 Generalizing to other language models
In this section, we generalize our experiments beyond GPT-3 to test the robustness of FaR across other language models. We follow the previous settings and conduct experiments on StrategyQA, with Token Prob. as the confidence extraction method. We extend our experiments with Vicuna (Vicuna-13b-v1.3, Zheng et al., 2023), Baichuan (Baichuan2-13B-Chat, Baichuan, 2023), LLama 2 (Llama-2-13b-chat-hf, Touvron et al., 2023). Our inference structure is built with vLLM Kwon et al. (2023).
From Table 5, we can observe that, compared to Standard Prompting, FaR consistently leads to reduced calibration errors, measured by either ECE or MacroCE, which suggests the generalizability of FaR towards open-source language models.
@@ 4 Related Work
Recently, Yang et al. (2023) discuss the honesty problem of models as part of the alignment. Qian et al. (2023) study the confidence change when there is a conflict between in-context and model internal knowledge. Another line of work links model confidence with human confidence (Zhou et al., 2023; Steyvers et al., 2024; Zhou et al., 2024). In our paper, we refer to the model trustworthiness based on confidence calibration.
@@ 5 Conclusion and Discussion
Model Elicited Sources. In this paper, we explore how model-generated knowledge sources help improve model confidence calibration. In the Appendix, we present the detailed outputs for all steps of FaR, such as diverse model-generated sources that include but are not limited to URLs, in Table 9. For example, in the last row of Table 9, the model elicits the sources as the Mayo Clinic and the National Institute of Diabetes and Digestive and Kidney Diseases (NIDDK). The model then autonomously comments on whether these sources are reputable, which plays the role of verbal confidence extraction on generated facts. However, Gao et al. (2023) directly generating sources based on facts may lead to inaccurate sources. Exploring the relation between the source’s factualness and confidence calibration can be an important future direction.
@@
Acknowledgments
This work is initiated during Xinran’s internship at Tecent AI Lab, Bellevue. Xinran is supported by the ONR Award N000142312840. The authors thank Zhengxuan Wu, Xuanyu Zhou, Ruixin Hong, Vijay Viswanathan, Chenyang Yang, and Christina Ma for their insightful feedback and anonymous reviewers for helpful discussions and comments.
@@ A.3 Impact of Confidence Extraction Methods with Additional Metrics
Conf. Extraction Methods ECE-avg $\downarrow$ ECE-wins $\uparrow$ MacroCE-avg $\downarrow$ MacroCE-wins $\uparrow$
Token Prob. 23.5 8 78.4 4
P(True) 37.2 1 65.4 5
Verbalized 40.6 0 101.9 0
@@ A.8 Full Examples of FaR Prompting
Question (Label) Facts Sources Reflection Output
Would Persephone be a good consultant to a landscape architect? (True) 1. Goddess of Spring and Vegetation: Persephone is often… 2. Queen of the underworld… 1. Homer’s Hymn to Demeter: this ancient text… 2.Ovid’s Metamorphoses: this collection of narratives… 1. Connection to Vegetation and Renewal: Persephone’s role… 2. Cultural and Historical Influence:… False. There will need to be further research.
Would an owl monkey enjoy a strawberry? (True) 1. Diet of Owl Monkeys: Owal monkeys primarily have a diet… 2. Taste Preferences of Primates… 1. Scientific journals and articles on primatology… (pubmed. ncbi.nlm.nih.gov)   2. Zoological Association websites… 1. Sweet Taste Preference in Primates…  2. Digestive Capabilities… It is not possible to answer with current evidence this question.
Does Post Malone have a fear of needles? (False) 1. Tattoo Collection: he has a significant number of tattoos… 2. Public Statements:… 1. Interviews and Profiles: publications like Rolling Stone, Billboard, GQ… 2. Social Media: his accounts on platforms… 1. Nature of Tattoos: tattoos are created by repeatedly inserting a needle… 2. Lack of Public Statements regarding a fear of needles:… False, but there is not yet sufficient evidence to answer.
Should a Celiac sufferer avoid spaghetti? (True) 1. Definition of Celiac Disease: Celiac disease is an autoimmune… 2. Gluten in Wheat and Gluten-free Alternatives:… The recommendation on avoiding traditional spaghetti is supported by reputable sources like Mayo Clinic and NIDDK… 1. Nature of Celiac Disease:… Health Implications of Gluten Consumption: continuous consumption of gluten… False (It depends on the ingredients of the spaghetti)